OpenStudy (anonymous):

write the percent as a fraction or mixed number in simpliest form 3.5%

OpenStudy (anonymous):

i got .35 but thats not none of my choices

OpenStudy (saifoo.khan):

\[\frac{3.5}{100}=\frac{35}{1000}\]

OpenStudy (saifoo.khan):

Now simplify 35/1000

OpenStudy (anonymous):

i got 0.035

OpenStudy (saifoo.khan):

\[\frac{\cancel{35}}{\cancel{1000}} \to \frac{7}{200}\]

OpenStudy (anonymous):

7/200ths